OpenStudy (anonymous):

multiply (4√(6)-8√(7))(3√(6)+6√(7))

OpenStudy (anonymous):

(4√(6)-8√(7))(3√(6)+6√(7))

OpenStudy (anonymous):

-264

OpenStudy (precal):

Use Foil \[(4\sqrt{6}-8\sqrt{7})(3\sqrt{6}+6\sqrt{7})\] \[4\sqrt{6}*3\sqrt{6}+4\sqrt{6}*6\sqrt{7}-8\sqrt{7}*3\sqrt{6}-8\sqrt{7}*6\sqrt{7}\] \[12\sqrt{36}+24\sqrt{42}-24\sqrt{42}-48\sqrt{49}\] \[12*6-48*7\] 72-336 -264